# The cone conjecture for primitive symplectic varieties over a field of characteristic zero and an application

### v1

@@ 3.1. Preliminaries
then we are done. Let $E$ be such a prime exceptional divisor. By our assumptions, we already know that the inequality is true if the orbit $I$ of $E$ for the action of the absolute Galois group is such that $W_{I}$ is finite, so let us assume that $W_{I}$ is infinite. By contradiction, suppose that $q_{\overline{X}}(E,\alpha)<0$. Since $\mathop{\rm Pos}\nolimits(X)$ is connected, there exists then $\lambda\in\mathop{\rm Pos}\nolimits(X)$ such that $q_{\overline{X}}(E,\lambda)=0$. As $\lambda$ is invariant under the action of the absolute Galois group, $\lambda$ is $q_{\overline{X}}$-orthogonal to every element in $I$ and is therefore fixed by the action of $W_{I}$. Consequently, $\lambda$ is also $q_{\overline{X}}$-orthogonal to $w(E)$ for any $w\in W_{I}$. Note that orbit $I$ is finite because $\mathop{\rm Pic}\nolimits(\overline{X})$ is of finite type so that the action of the Galois group $G_{F}$ factors through a finite quotient. By [undefaar, Theorem 1] and noting that $wr_{E}w^{-1}=r_{w(E)}$, we deduce that $W_{I}\cdot E=\left\{w(E)\penalty 10000\ \middle|\penalty 10000\ w\in W_{I}\right\}$ is an infinite set. Denote by $N$ the set of all the classes $\beta\in\mathop{\rm Pic}\nolimits(\overline{X})_{\mathbb{R}}$ such that $q_{\overline{X}}(\beta,\beta)<0$. From what we have just said, $\lambda$ is $q_{\overline{X}}$-orthogonal to an infinite number of elements of $N$. However, arguing as in [undefak, Remark 8.2.3], we obtain that the following union
@@
is locally finite, and we have reached a contradiction. ∎
@@
Let $F$ be a field of characteristic $0$ and let $X$ be a primitive symplectic variety over $F$ with $\mathbb{Q}$-factorial terminal singularities and $b_{2}(X)\geq 5$. Then $\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ is a fundamental domain for the action of $R_{X}$ on $\mathop{\rm Pos}\nolimits(X)$.
@@
We denote by $\overline{X}$ the base change $X_{\overline{F}}$, and by $G_{F}$ the absolute Galois group of $F$. First, we show that for every $\alpha\in\mathop{\rm Pos}\nolimits(X)$, there exists $g\in R_{X}$ such that $g\alpha\in\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$. If the stabilizer of $\alpha$ for the action of $W_{\overline{X}}$ on $\mathop{\rm Pic}\nolimits(\overline{X})_{\mathbb{R}}$ is trivial, there is then a unique $g\in W_{\overline{X}}$ such that $g\alpha\in\mathop{\overline{\rm Mov}}\nolimits(\overline{X})\cap\mathop{\rm Pos}\nolimits(\overline{X})$ by Proposition 3.14. We claim that $g\in R_{X}$. Indeed, for any $\sigma\in G_{F}$, we have
@@
$x\in\bigcup_{E\in{\mathcal{P}}_{\overline{X}}}E^{\perp}\cap\mathop{\rm Pos}\nolimits(\overline{X}),$ (3.6)
where ${\mathcal{P}}_{\overline{X}}$ is the set of prime exceptional divisors of $\overline{X}$. Arguing as in [undefak, Remark 8.2.3], the above union is locally finite. Since $X$ admits an ample divisor, $\mathop{\rm Pic}\nolimits(X)$ is not contained in any of the walls appearing in (3.6). This allows to construct a sequence $(\alpha_{n})_{n\geq 0}$ of elements of $\mathop{\rm Pos}\nolimits(X)$ tending to $\alpha$, and all lying in the interior of the same chamber of $\mathop{\rm Pos}\nolimits(\overline{X})$ cut out by the walls above. From what we have said before, we deduce that there is a unique $g\in R_{X}$ such that $g\alpha_{n}\in\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ for every $n\geq 0$. By continuity and passing to the limit, we obtain $g\alpha\in\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$.
Next, we show that for all non-trivial element $g\in R_{X}$, $\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ and $g\cdot\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ intersect only at their boundaries. Suppose that $x\in\mathop{\rm Pos}\nolimits(X)$ lies in $\mathop{\overline{\rm Mov}}\nolimits(X)\cap g\cdot\mathop{\overline{\rm Mov}}\nolimits(X)$ for some non-trivial element $g\in R_{X}$. By Proposition 3.14, $x$ lies in the boundary of $\mathop{\overline{\rm Mov}}\nolimits(\overline{X})$, and Lemma 3.5 shows that $x$ lies in the boundary of $\mathop{\overline{\rm Mov}}\nolimits(X)$, which concludes this proof. ∎
@@ 3.2. The absolute movable cone conjecture
Before proving Theorem A.(1), we recall a few facts about the group $\mathop{\rm Bir}\nolimits(X)$ of a projective variety $X$ over an algebraically closed field $L$ of characteristic $0$. For a more detailed account, please refer to [undefh, Section 3].
Definition 3.21 ([undefaf] - §1-1.1).
Let $X$ be an algebraic variety over an algebraically closed field $L$. A birational automorphism of $X$ is defined to be a birational map from $X$ to itself. Given a rational map $f:X\dashrightarrow X$, $f$ is a birational automorphism if and only if the graph $\Gamma_{f}\subset X\times X$ is a birational correspondence, i.e. each of the two projections of $X\times X$, when restricted to $\Gamma_{f}$, are birational morphisms.
Definition 3.22 (Definition 2.1 - [undefaf]).
Let $X$ be a projective variety over $L$ and let $S$ be a locally Noetherian scheme over $L$. A flat family of $S$-birational automorphisms of $X$ is a closed subscheme $Z\subset S\times X\times X$, flat over $S$ and such that for all points $t\in S$, the fibres $Z_{t}$ over $t$ are birational correspondences.
Proposition 3.23 (Proposition 1.7 - [undefaf]).
Let $X$ be a projective variety over an algebraically closed field $L$ of characteristic $0$. Denote by $\operatorname{(locNoethSch/L)}$ the full subcategory of $\operatorname{(Sch/L)}$ consisting of locally Noetherian schemes over $L$. The contravariant functor
$\mathop{\rm Bir}\nolimits^{\text{flat}}_{X/L}:\operatorname{(locNoethSch/L)}\to(\text{Set}),\penalty 10000\ \penalty 10000\ S\mapsto\left\{\text{flat families of $\displaystyle S$-birational automorphisms of $\displaystyle X$}\right\}$
is representable by a scheme $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(X)$.
@@
Theorem 3.24 (Theorem 2.1 - [undefag]).
Let $X$ be a non-uniruled projective variety over an algebraically closed field $L$ of characteristic $0$. The following holds.
@@
First, we show that $X$ is non-uniruled. By contradiction, suppose that there exists a variety $Y$ over $L$ and a dominant rational map $\phi:Y\times\mathbb{P}^{1}_{L}\dashrightarrow X$. Take a finitely generated subfield $L^{\prime}\subset L$ such that $Y$, $\mathbb{P}^{1}_{L}$, $X$ and $\phi$ are defined over $L^{\prime}$. After fixing an embedding $\overline{L^{\prime}}\subset\mathbb{C}$, we obtain that $\phi_{\mathbb{C}}:Y_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}\dashrightarrow X_{\mathbb{C}}$ is a dominant rational map, which means that $X_{\mathbb{C}}$ is uniruled. Note that $X_{\mathbb{C}}$ is a projective primitive symplectic variety by Lemma 2.3 and Lemma 2.4. Let $\pi:\tilde{X}\to X_{\mathbb{C}}$ be a resolution of singularities. $\tilde{X}$ is uniruled since $X_{\mathbb{C}}$ is, therefore $h^{0}(\tilde{X},\mathcal{O}_{\tilde{X}}(K_{\tilde{X}}))=0$, but this contradicts the fact that the pullback of the symplectic $2$-form on $X_{\mathbb{C}}$ extends to a regular $2$-form on $\tilde{X}$. Thus, we deduce that $X$ is non-uniruled.
@@
Next, we apply Theorem 3.24 to $X$. Given that $h^{1}(X,\mathcal{O}_{X})=0$ and that $X$ has rational singularities by Lemma 2.7 and [undefv, Théorème 1], it follows that the dimension of $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(X)$ is $0$. Let $Y$ be a smooth projective variety birationally equivalent to $X$ as in Theorem 3.24. Note that $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)_{\text{red}}$ is smooth over $L$ by [undefr, Exposé VI.B, Corollaire 1.6.1]. Since $\dim\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(X)=\dim\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)$ by [undefaf, Corollary 2.8], the group scheme $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)_{\text{red}}$ is of dimension $0$ and is therefore étale over $L$, equivalently it is a constant group scheme. Consequently, the morphism $\mathop{\rm Spec}\nolimits M\to\mathop{\rm Spec}\nolimits L$ induces an isomorphism between $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)_{\text{red}}(M)=\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)(M)$ and $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)_{\text{red}}(L)=\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)(L)$, which are respectively isomorphic to $\mathop{\rm Bir}\nolimits(Y_{M})$ and $\mathop{\rm Bir}\nolimits(Y)$. Since $Y$ is birationally equivalent to $X$, these two groups are respectively isomorphic to $\mathop{\rm Bir}\nolimits(X_{M})$ and $\mathop{\rm Bir}\nolimits(X)$, and this completes the proof. ∎
@@
Recall that if $X$ and $X^{\prime}$ are terminal $\mathbb{Q}$-factorial primitive symplectic varieties over a field $L$ of characteristic $0$, then a birational map $\phi:X\dashrightarrow X^{\prime}$ is an isomorphism in codimension $1$ by [undefap, Corollary 3.54]. Therefore, if $X^{\prime}=X$, then $\phi$ induces an isometry $\phi^{*}\in O(\mathop{\rm Pic}\nolimits(X))$. In order to prove the birational cone conjecture, we need to show that $\mathop{\rm Bir}\nolimits(X_{\overline{F}})$ is a finitely generated group. The strategy consists in adapting [undefi, Lemma 2 and Theorem 2] to the singular case, relying on the results of [undefat] to generalize [undefaas, Lemma 2.1 and Proposition 2.2].
@@
For a complex projective primitive symplectic variety $X$, the group of locally trivial monodromy operators preserving the pure Hodge structure on $H^{2}(X,\mathbb{Z})_{\mathop{\rm tf}\nolimits}$ as in [undefat, Definition 2.16] will be denoted by $\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits(X)$ .
Theorem 3.27 (Theorem 4.2 - [undefat]).
@@
We adapt the arguments of [undefi, Lemma 2]. Denote by $T(X)\subset H^{2}(X,\mathbb{Z})$ the transcendental “lattice” of $X$, i.e. the smallest integral Hodge structure such that $T(X)_{\mathbb{C}}:=T(X)\otimes_{\mathbb{Z}}\mathbb{C}$ contains $H^{2,0}(X)$. By [undefaap, Theorem 8] and arguing as in [undefak, Lemma 3.1], we have $T(X)_{\mathop{\rm tf}\nolimits}=\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}^{\perp}$. Set $T(X)_{\mathbb{R}}:=T(X)\otimes_{\mathbb{Z}}\mathbb{R}$ and $E:=\left(H^{2,0}(X)\oplus H^{0,2}(X)\right)\cap H^{2}(X,\mathbb{R})$. Denote by $\tau$ the rank of $T(X)$. By definition of the transcendental lattice and by Definition/Proposition 2.13, $(T(X)_{\mathop{\rm tf}\nolimits},q_{X})$ is non-degenerate and is of signature $(2,\tau-2)$. Furthermore, we have an orthogonal decomposition
@@
Let $L$ be an algebraically closed field of characteristic $0$ and $X$ a terminal $\mathbb{Q}$-factorial primitive symplectic variety over $L$ with $b_{2}(X)\geq 5$. Then the birational group $\mathop{\rm Bir}\nolimits(X)$ of $X$ is finitely generated.
@@
Let $L_{0}\subset L$ be a finitely generated subfield such that $X$ is defined over $L_{0}$ and denote by $X_{0}$ (respectively $X_{0,\mathbb{C}}$) the corresponding primitive symplectic variety over $L_{0}$ (respectively the base change of $X_{0}$ to $\mathbb{C}$). From Lemma 3.25, we know that $\mathop{\rm Bir}\nolimits(X)\simeq\mathop{\rm Bir}\nolimits(X_{0,\mathbb{C}})$, so we can assume that $L=\mathbb{C}$. Thanks to Lemma 3.28, in order to show that $\mathop{\rm Bir}\nolimits(X)$ is finitely generated, it is sufficient to show that $\mathop{\rm Bir}\nolimits^{0}(X)$ is finitely generated. Consider the restriction morphism $\rho:\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits\to O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ and define:
${\mathcal{C}}^{\#}:=\mathop{\rm Pos}\nolimits(X)\cap\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits},\penalty 10000\ \penalty 10000\ O^{\#}(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}):=\left\{g\in O(\mathop{\rm Pic}\nolimits(X))\penalty 10000\ \middle|\penalty 10000\ g({\mathcal{C}}^{\#})={\mathcal{C}}^{\#}\right\}.$
By [undefat, Lemma 6.3.(1)], the image $\Gamma$ of $\rho$ is a finite index subgroup of $O^{\#}(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ and is in particular finitely generated by [undefj, Theorem 6.12]. Using the notation from the proof of Lemma 3.28, define $\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits(X)_{T}:=\left\{\phi\in\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits(X)\penalty 10000\ \middle|\penalty 10000\ \phi|_{T(X)}=\text{id}\right\}$ and $\Gamma_{T}=\rho(\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits(X)_{T})\subset\Gamma$. The group $\Gamma_{T}$ is an arithmetic subgroup of $\Gamma$, so is finitely generated. Let ${\mathcal{C}}{\mathcal{P}}_{X}\subset\mathop{\rm Pic}\nolimits(X)_{\mathbb{Q}}$ be the set of classes of prime exceptional divisors of $X$ and define

### v2

@@ 3.1. Preliminaries
then we are done. Let $E$ be such a prime exceptional divisor. By our assumptions, we already know that the inequality is true if the orbit $I$ of $E$ under the action of the absolute Galois group is such that $W_{I}$ is finite, so let us assume that $W_{I}$ is infinite. By contradiction, suppose that $q_{\overline{X}}(E,\alpha)<0$. Since $\mathop{\rm Pos}\nolimits(X)$ is connected, there exists $\lambda\in\mathop{\rm Pos}\nolimits(X)$ such that $q_{\overline{X}}(E,\lambda)=0$. As $\lambda$ is invariant under the action of the absolute Galois group, $\lambda$ is $q_{\overline{X}}$-orthogonal to every element in $I$ and is therefore fixed by the action of $W_{I}$. Consequently, $\lambda$ is also $q_{\overline{X}}$-orthogonal to $w(E)$ for any $w\in W_{I}$. Note that the orbit $I$ is finite because $\mathop{\rm Pic}\nolimits(\overline{X})$ is of finite type, so that the action of the Galois group $G_{F}$ factors through a finite quotient. By [undefaar, Theorem 1] and noting that $wr_{E}w^{-1}=r_{w(E)}$, we deduce that $W_{I}\cdot E=\left\{w(E)\penalty 10000\ \middle|\penalty 10000\ w\in W_{I}\right\}$ is an infinite set. Denote by $N$ the set of all the classes $\beta\in\mathop{\rm Pic}\nolimits(\overline{X})_{\mathbb{R}}$ such that $q_{\overline{X}}(\beta,\beta)<0$. From what we have just said, $\lambda$ is $q_{\overline{X}}$-orthogonal to an infinite number of elements of $N$. However, arguing as in [undefaj, Remark 8.2.3], we obtain that the following union
@@
is locally finite, which yields a contradiction. ∎
@@
Let $F$ be a field of characteristic $0$ and let $X$ be a projective primitive symplectic variety over $F$ with $b_{2}(X)\geq 5$ such that $\overline{X}:=X_{\overline{F}}$ has $\mathbb{Q}$-factorial terminal singularities. Then $\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ is a fundamental domain under the action of $R_{X}$ on $\mathop{\rm Pos}\nolimits(X)$.
@@
We denote by $\overline{X}$ the base change $X_{\overline{F}}$, and by $G_{F}$ the absolute Galois group of $F$. First, we show that for every $\alpha\in\mathop{\rm Pos}\nolimits(X)$, there exists $g\in R_{X}$ such that $g\alpha\in\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$.
If the stabilizer of $\alpha$ under the action of $W_{\overline{X}}$ on $\mathop{\rm Pic}\nolimits(\overline{X})_{\mathbb{R}}$ is trivial, there is a unique $g\in W_{\overline{X}}$ such that $g\alpha\in\mathop{\overline{\rm Mov}}\nolimits(\overline{X})\cap\mathop{\rm Pos}\nolimits(\overline{X})$ by Proposition 3.14. We claim that $g\in R_{X}$. Indeed, for any $\sigma\in G_{F}$, we have
@@
If now $\alpha$ has a non-trivial stabilizer, we must have by Proposition 3.14 that
$\alpha\in\bigcup_{E\in{\mathcal{P}}_{\overline{X}}}E^{\perp}\cap\mathop{\rm Pos}\nolimits(\overline{X}),$ (3.6)
where ${\mathcal{P}}_{\overline{X}}$ is the set of prime exceptional divisors of $\overline{X}$. Arguing as in [undefaj, Remark 8.2.3], the above union is locally finite. Since $X$ admits an ample divisor, $\mathop{\rm Pic}\nolimits(X)$ is not contained in any of the walls appearing in (3.6). This allows to construct a sequence $(\alpha_{n})_{n\geq 0}$ of elements of $\mathop{\rm Pos}\nolimits(X)$ tending to $\alpha$, and all lying in the interior of the same chamber of $\mathop{\rm Pos}\nolimits(\overline{X})$ cut out by the walls above. From what we have said before, we deduce that there is a unique $g\in R_{X}$ such that $g\alpha_{n}\in\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ for every $n\geq 0$. By continuity, it follows that $g\alpha\in\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$.
Next, we show that for all non-trivial element $g\in R_{X}$, $\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ and $g\cdot\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ intersect only at their boundaries. Suppose that $\alpha\in\mathop{\rm Pos}\nolimits(X)$ lies in $\mathop{\overline{\rm Mov}}\nolimits(X)\cap g\cdot\mathop{\overline{\rm Mov}}\nolimits(X)$ for some non-trivial element $g\in R_{X}$. By Proposition 3.14, $\alpha$ lies in the boundary of $\mathop{\overline{\rm Mov}}\nolimits(\overline{X})$, and Lemma 3.5 shows that $\alpha$ lies in the boundary of $\mathop{\overline{\rm Mov}}\nolimits(X)$, which concludes this proof. ∎
@@ 3.2. The absolute movable cone conjecture
Before proving Theorem A.(1), we recall a few facts about the group $\mathop{\rm Bir}\nolimits(X)$ of a projective variety $X$ over an algebraically closed field $L$ of characteristic zero. We denote by $\operatorname{(LNSch/L)}$ the full subcategory of $\operatorname{(Sch/L)}$ consisting of locally Noetherian schemes over $L$. In the following, we adopt the notations of [undefg, Section 3].
Definition 3.21.
Let $X$ be a variety over an algebraically closed field $L$ of characteristic zero. We define the following functor.
[…]
Lemma 3.22 (Lemma 3.3 - [undefg]).
Let $X$ be a variety over an algebraically closed field $L$ of characteristic zero. For every locally Noetherian scheme $S$ over $L$, we have a natural bijection
[…]
which sends $f\in\mathop{\rm Bir}\nolimits_{X/L}(S)$ to $\Gamma_{f}:=\overline{\left\{(s,x,q(f(s,x))\penalty 10000\ \middle|\penalty 10000\ (s,x)\in\operatorname{dom}(f)\right\}}$, where $q:S\times X\to X$ is the projection.
Definition 3.23 (Definition 2.1 - [undefae]).
Let $X$ be a projective variety over an algebraically closed field $L$ of characteristic zero, and $S$ a locally Noetherian scheme over $L$. A flat family of birational transformations of $X$ over $S$ is a closed subscheme $Z\subset S\times X\times X$ which is flat over $S$, and such that for every point $s\in S$, the fibre $Z_{s}$ is the graph of an element of $\mathop{\rm Bir}\nolimits_{X}(k(s))$.
Proposition 3.24 (Proposition 1.7 - [undefae]).
Let $X$ be a projective variety over an algebraically closed field $L$ of characteristic $0$. The functor
$\mathop{\rm Bir}\nolimits^{\text{flat}}_{X/L}:\operatorname{(LNSch/L)}\to(\text{Set}),\penalty 10000\ \penalty 10000\ S\mapsto\left\{f\in\mathop{\rm Bir}\nolimits_{X/L}(S)\penalty 10000\ \middle|\penalty 10000\ \Gamma_{f}\text{ is a flat family of birational transformations}\right\}$
is representable by a scheme $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(X)$, where $\Gamma_{f}$ is the graph of $f$ as defined in Lemma 3.22.
@@
Theorem 3.25 (Theorem 2.1 - [undefaf]).
Let $X$ be a non-uniruled projective variety over an algebraically closed field $L$ of characteristic zero. The following holds.
@@
First, we show that $X$ is non-uniruled. Suppose by contradiction that there exist a variety $Y$ over $L$ and a dominant rational map $\phi:Y\times\mathbb{P}^{1}_{L}\dashrightarrow X$. Take a finitely generated subfield $L^{\prime}\subset L$ such that $Y$, $\mathbb{P}^{1}_{L}$, $X$ and $\phi$ are defined over $L^{\prime}$, and fix an embedding $\overline{L^{\prime}}\subset\mathbb{C}$. The base change $\phi_{\mathbb{C}}:Y_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}\dashrightarrow X_{\mathbb{C}}$ is a dominant rational map, which means that $X_{\mathbb{C}}$ is uniruled. Note that $X_{\mathbb{C}}$ is a projective primitive symplectic variety by Lemma 2.3 and Lemma 2.4. Let $\pi:\tilde{X}\to X_{\mathbb{C}}$ be a resolution of singularities. Since $X_{\mathbb{C}}$ is uniruled, $\tilde{X}$ is also uniruled, and therefore $h^{0}(\tilde{X},\mathcal{O}_{\tilde{X}}(K_{\tilde{X}}))=0$. This contradicts the fact that the pullback of the non-zero symplectic $2$-form on $X_{\mathbb{C}}$ extends to a regular $2$-form on $\tilde{X}$. Hence, $X$ is non-uniruled.
@@
Next, we apply Theorem 3.25 to $X$. Given that $h^{1}(X,\mathcal{O}_{X})=0$ and that $X$ has rational singularities by Lemma 2.7 and [undefu, Théorème 1], it follows that the dimension of $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(X)$ is zero. Let $Y$ be a smooth projective variety birationally equivalent to $X$ as in Theorem 3.25. By [undefq, Exposé VI.B, Corollaire 1.6.1], $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)_{\text{red}}$ is smooth over $L$. Since $\dim\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(X)=\dim\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)$ by [undefae, Corollary 2.8], the group scheme $\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)_{\text{red}}$ is of dimension $0$ and is therefore étale over $L$, equivalently it is a constant group scheme. It follows that the morphism $\mathop{\rm Spec}\nolimits M\to\mathop{\rm Spec}\nolimits L$ induces an isomorphism
$\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)_{\text{red}}(M)\simeq\mathop{\rm\textbf{Bir}}\nolimits_{L}^{\textbf{flat}}(Y)_{\text{red}}(L),$
which simplifies to an isomorphism $\mathop{\rm Bir}\nolimits(Y_{M})\simeq\mathop{\rm Bir}\nolimits(Y)$ by definition of $\mathop{\rm Bir}\nolimits^{\text{flat}}_{Y/L}$. Since $Y$ is birationally equivalent to $X$, these two groups are respectively isomorphic to $\mathop{\rm Bir}\nolimits(X_{M})$ and $\mathop{\rm Bir}\nolimits(X)$, and this completes the proof. ∎
@@
Recall that if $X$ and $X^{\prime}$ are projective primitive symplectic varieties over a field $L$ of characteristic zero with $\mathbb{Q}$-factorial terminal singularities, then a birational map $\phi:X\dashrightarrow X^{\prime}$ is an isomorphism in codimension one by [undefao, Corollary 3.54]. Therefore, if $X^{\prime}=X$, then $\phi$ induces an isometry $\phi^{*}\in O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$. In order to prove the birational cone conjecture, we need to show that $\mathop{\rm Bir}\nolimits(X_{\overline{F}})$ is a finitely generated group. The strategy consists in adapting [undefh, Lemma 2 and Theorem 2] to the singular case, relying on the results of [undefas] to generalize [undefaas, Lemma 2.1 and Proposition 2.2].
@@
For a complex projective primitive symplectic variety $X$, the group of locally trivial monodromy operators preserving the pure Hodge structure on $H^{2}(X,\mathbb{Z})_{\mathop{\rm tf}\nolimits}$ as in [undefas, Definition 2.16] will be denoted by $\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits(X)$ .
Theorem 3.28 (Theorem 4.2 - [undefas]).
@@
We adapt the arguments of [undefh, Lemma 2]. Denote by $T(X)\subset H^{2}(X,\mathbb{Z})$ the transcendental “lattice” of $X$, i.e. the smallest integral Hodge structure such that $T(X)_{\mathbb{C}}:=T(X)\otimes_{\mathbb{Z}}\mathbb{C}$ contains $H^{2,0}(X)$. By [undefaap, Theorem 8] and arguing as in [undefaj, Lemma 3.1], we have $T(X)_{\mathop{\rm tf}\nolimits}=\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}^{\perp}$. Set $T(X)_{\mathbb{R}}:=T(X)\otimes_{\mathbb{Z}}\mathbb{R}$ and $E:=\left(H^{2,0}(X)\oplus H^{0,2}(X)\right)\cap H^{2}(X,\mathbb{R})$. Denote by $\tau$ the rank of $T(X)$. By definition of the transcendental lattice and by Definition/Proposition 2.13, $(T(X)_{\mathop{\rm tf}\nolimits},q_{X})$ is non-degenerate and is of signature $(2,\tau-2)$. Furthermore, we have an orthogonal decomposition
@@
Let $L$ be an algebraically closed field of characteristic $0$ and $X$ a projective primitive symplectic variety over $L$ with $\mathbb{Q}$-factorial terminal singularities and $b_{2}(X)\geq 5$. The birational group $\mathop{\rm Bir}\nolimits(X)$ of $X$ is finitely generated.
@@
Let $L_{0}\subset L$ be a finitely generated subfield such that $X$ is defined over $L_{0}$ and denote by $X_{0}$ (respectively $X_{0,\mathbb{C}}$) the corresponding primitive symplectic variety over $L_{0}$ (respectively the base change of $X_{0}$ to $\mathbb{C}$). From Lemma 3.26, we know that $\mathop{\rm Bir}\nolimits(X)\simeq\mathop{\rm Bir}\nolimits(X_{0,\mathbb{C}})$, so we can assume that $L=\mathbb{C}$. Thanks to Lemma 3.29, in order to show that $\mathop{\rm Bir}\nolimits(X)$ is finitely generated, it is sufficient to show that $\mathop{\rm Bir}\nolimits^{0}(X)$ is finitely generated. Consider the restriction morphism $\rho:\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits\to O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ as well as
${\mathcal{C}}^{\#}:=\mathop{\rm Pos}\nolimits(X)\cap\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits},\penalty 10000\ \text{ and }\penalty 10000\ O^{\#}(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}):=\left\{g\in O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})\penalty 10000\ \middle|\penalty 10000\ g({\mathcal{C}}^{\#})={\mathcal{C}}^{\#}\right\}.$
By [undefas, Lemma 6.3.(1)], the image $\Gamma$ of $\rho$ is a finite index subgroup of $O^{\#}(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ and is in particular finitely generated by [undefi, Theorem 6.12]. Using the notation from the proof of Lemma 3.29, define $\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits(X)_{T}:=\left\{\phi\in\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits(X)\penalty 10000\ \middle|\penalty 10000\ \phi|_{T(X)}=\text{id}\right\}$ and $\Gamma_{T}=\rho(\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits(X)_{T})\subset\Gamma$. The group $\Gamma_{T}$ is an arithmetic subgroup of $\Gamma$, so is finitely generated. Let ${\mathcal{C}}{\mathcal{P}}_{X}\subset\mathop{\rm Pic}\nolimits(X)_{\mathbb{Q}}$ be the set of classes of prime exceptional divisors of $X$ and define